OpenStudy (anonymous):

[2.06] Solve the equation x-5/3 = 6 for x. 13 14 23 10

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large{x - 5 \over 3} = 6 }\) Is that your question?

OpenStudy (venomblast):

the answer is 23

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large{x - 5 \over \color{Blue}{\cancel3}} \times \cancel 3 = 6 \times 3 }\) \( \color{Black}{\Rightarrow x - 5 = 18 }\) \( \color{Black}{\Rightarrow x = 23 }\)

OpenStudy (venomblast):

first get rid of the fraction. multiply by 3 both side. so you will get x-5=18 then add 5 and you will get your answer

Parth (parthkohli):

You must multiply 3 to both sides, it'd be cancelled from the denominator. Then add 5 to both sides.